OpenStudy (latavia.2k18):

If f(x)=3x-8 and g(x)=4-7x, what is f(g(x))?

OpenStudy (jaedyn):

https://youtu.be/Qu-c-XWLNf0

OpenStudy (jaedyn):

subscribe

OpenStudy (3mar):

May I help?

OpenStudy (latavia.2k18):

Yes

OpenStudy (3mar):

Thanks.

OpenStudy (3mar):

This form of operations on function is called: function of the function. In other simple words: f(g(x)) means f is a function of the function g(x) which means: replace every x in the function f(x) with g(x). For our case: f(x)=3x-8 g(x)=4-7x so that f(g(x))=3[g(x)]-8=3[4-7x]-8=12-21x-8=4-21x That is it! I hope that helps!

OpenStudy (3mar):

@Latavia.2k18

OpenStudy (latavia.2k18):

Omg thank you this really helped me a lot @3mar

OpenStudy (3mar):

Thank you for learning! Any Help... Any Time...

OpenStudy (3mar):

Any more questions?